Skip to content
Past Exams

Mathematics Final Exam With Answers (10 Solved Questions)

Mathematics Final Exam With Answers improves your recall with a structured past exam review.

Preview 3 of 7 pages

See the first 3 pages exactly as they appear in the document.

  1. Page 1
    Read page 1 as text

    Portage Learning MATH 110 Introduction to Statistics Final Exam Question 1 Not yet graded / 10 pts You may find the following files helpful throughout the exam: The following pie chart shows the percentages of total items sold in a month in a certain fast food restaurant. A total of 4700 fast food items were sold during the month. a.) How many were fish? b.) How many were french fries? Your Answer: a. fish 4700(0.28)=1316 b. French fries 4700(0.4)=1880 a.) Fish : 4700(.28) = 1316 b.) French Fries: 4700(.40) = 1880 Question 2 Consider the following data: 430 389 414 401 466 421 399 387 450 407 392 410 440 417 471 Find the 40th percentile of this data.

  2. Page 2
  3. Page 3
    Read page 3 as text

    There are a total of fifteen numbers, so n= 15. In order to find the percentiles, we must put the numbers in ascending order: 387 389 392 399 401 407 410 414 417 421 430 440 450 466 471 For the 40th percentile: Therefore, the 40th percentile index for this data set is the 6th observation. In the list above, the 6th observation is 407. Question 3 In a tri - state conference, 60% attendees are from California, 25% from Oregon, and 15% from Washington. As it turns out 6 % of the attendees from California, 17% of the attendees from Oregon, and 12% of the attendees from Washington came to the conference by train. If an attendee is selected at random and found to have arrived by train, what is the probability that the person is from Washington? P(Train│C)=.06.. P(Train│O)=.17.. P(Train│W)=.12.. P(C)=.60,P(O)=.25,P(W)=.15. We want to find P(W│Train), so use: Question 4 Find each of the following probabilities: a. Find P(Z ≤ - 0.87) . b. Find P(Z ≥ .93) . c. Find P( - .59 ≤ Z ≤ - .36). a. P(Z ≤ - 0.87)= .19215. b. P(Z ≥ .93)=1 - .82381= .17619.

4 more pages in the full document

Unlock it once and it stays in your library, ready to chat with or turn into flashcards.

Unlock the full document
See all

Study this document with CramX

Once it is in your library, every tool can work from it.